### v1

Abstract.
Let $G$ be a countable group that admits a minimal equicontinuous action on the Cantor set (i.e., an extension of an infinite residually finite group). We show that the family of almost automorphic actions of $G$ on the Cantor set, is a test for amenability for $G$. More precisely, we show that if $G$ is non-amenable, then for every minimal equicontinuous system $(Y,G)$, such that $Y$ is a Cantor set, there exists a Toeplitz subshift with no invariant probability measures, whose maximal equicontinuous factor is $(Y,G)$. Consequently, we obtain that $G$ is amenable if and only if every Toeplitz $G$-subshift has at least one invariant probability measure.
Key words and phrases:
amenability test, subshifts, almost automorphic, invariant measures.
2020 Mathematics Subject Classification:
Primary: 37B05, 37B10; Secondary: 20E26
J. Gómez was supported by ANID/ Doctorado Nacional No. 21200054
1. Introduction
A classical result states that non-amenable countable groups are characterized as those that admit continuous actions on compact metric spaces with no invariant probability measures (see [1, 17, 33] ). Motivated by this characterization and in response to a question from Grigorchuk, Giordano and de la Harpe show in [18] that to test the non-amenability of a countable group, it suffices to find such an action on the Cantor set. In other words, they show that the class of continuous actions of a countable group $G$ on the Cantor set is a test family for amenability for $G$. A direct application of [18] allows the restriction of the test family even further, for it is sufficient to consider only symbolic systems (see, for example, Lemma 10). Moreover, it is possible to narrow down the size of the alphabet on which the symbolic systems that form the test family for amenability are defined. For example, as a consequence of [27, Theorem 1.2], it can be deduced that for any non-amenable group $G$ there exists a $G$-subshift on $n=n(G)$ letters which admits no invariant probability measure. Recently, Frisch, Kechris, Shinko, and Vidnyánszky show in [14] that the non-amenability of a countable group can be tested within the class of subshifts defined on a two-letter alphabet. Thus, it is natural to wonder whether finding other test families for amenability is possible.
This article focuses on the family of almost automorphic systems, defined as the topological dynamical systems whose maximal equicontinuous factor is minimal and the associated factor map is almost one-to-one (see [15, 29, 30]). The symbolic, almost automorphic systems whose maximal equicontinuous factors are Cantor systems are precisely the Toeplitz $G$-subshifts. These subshifts exhibit a variety of exciting behaviors ([5, 10, 12, 19, 21, 28, 32]), such as the realization of all the metrizable Choquet simplex as spaces of invariant probability measures, and all the possible values for the topological entropy, in the case when the group $G$ is residually finite and amenable (see [7, 11, 24, 32]).
In this article, we show that the family of Toeplitz subshifts is a test for amenability for countable groups which are extensions of infinite residually finite groups. Additionally, this result gives an answer to the question posed by Veech in [30, Page 823] about the existence of almost automorphic systems without invariant measures, which was previously announced by Furstenberg and Weiss in [16].
Our first result is the following:
Theorem 1.
Let $(Y,G)$ be a minimal topological dynamical system, with $Y$ compact and metric. Let $(Z,G)$ be a minimal equicontinuous system, with $Z$ a Cantor set. Suppose there exists a factor map $\pi:Y\to Z$. Then there exist a minimal topological dynamical system $(X,G)$ and an almost one-to-one factor map $\tau:X\to Z$, such that
  1. (1)
    There exists a factor map $\psi:Y\to X$, such that $\psi$ is one-to-one on $\pi^{-1}(X_{1})$, where $X_{1}\subseteq Z$ is a full measure $G$-invariant set.
  2. (2)
    $\tau\circ\psi=\pi$.
  3. (3)
    The map $\psi$ induces an affine bijection between the spaces of invariant probability measures of $(Y,G)$ and $(X,G)$.
In particular, $(Y,G)$ has no invariant probability measures if and only if $(X,G)$ has no invariant probability measures.
Furthermore, if $(Y,G)$ is a Cantor minimal system, then so is $(X,G)$.
Theorem 1 was inspired by the techniques developed in [13], [16] and [31, Theorem 5.2]. It can be interpreted as a particular generalization of [13, Theorem 4.3] for group actions.
The following result allows us to get symbolic almost one-to-one extensions with no invariant measures.
Proposition 1.
Let $G$ be a non-amenable countable group. Let $(Y,G)$ be a minimal topological dynamical system, with $Y$ compact and metric, verifying the following properties:
  1. (1)
    There exists a minimal Cantor system $(X,G)$ without invariant probability measures, that is an almost one-to-one extension of $(Y,G)$.
  2. (2)
    There exists a minimal subshift $(Z,G)$ that is an almost one-to-one extension of $(Y,G)$.
Then there exists a minimal subshift $(\tilde{Z},G)$, with no invariant probability measures, that is an almost one-to-one extension of $(Y,G)$.
Combining Theorem 1 and Proposition 1, we get the following result:
Theorem 2.
Let $G$ be a non-amenable group. Then for every minimal equicontinuous system $(Y,G)$, such that $Y$ is a Cantor set, there exists a Toeplitz $G$-subshift with no invariant probability measures, whose maximal equicontinuous factor is $(Y,G)$.
As a corollary, we get the following characterization of amenability:
Corollary 1.
Let $G$ be a countable group which is an extension of an infinite residually finite group. The following statements are equivalent:
  1. (1)
    $G$ is amenable
  2. (2)
    Every almost automorphic topological dynamical system $(X,G)$ has invariant probability measures.
  3. (3)
    Every almost automorphic topological dynamical system $(X,G)$, whose maximal equicontinuous factor is a Cantor system, has invariant probability measures.
  4. (4)
    Every Toeplitz $G$-subshift has invariant probability measures.
  5. (5)
    There exists a minimal equicontinuous system $(Y,G)$, with $Y$ a Cantor set, such that every symbolic almost one-to-one extension of $(Y,G)$ has invariant probability measures.
This article is organized as follows: In Section 2, we recall concepts such as Toeplitz subshifts and odometers and their main associated results. In Section 3, we prove Theorem 1. Finally, in Section 4, we prove Proposition 1, and we deduce Theorem 2 and Corollary 1.
Acknowledgements
The authors warmly thank Raimundo Briceño, Paulina Cecchi, Sebastián Donoso, Tomasz Downarowicz, Patricio Pérez and Samuel Petite, for fruitful discussions.
2. Definitions and background.
A topological dynamical system is a pair $(X,G)$ where $X$ is a compact Hausdorff space and $G$ is a countable discrete group acting continuously on $X$. We call $X$ the phase space of the system. For every $x\in X$ and $g\in G$, we denote $gx$ the action of $g$ on $x$. The dynamical system $(X,G)$ is free if $gx=x$ implies $g=1_{G}$, for every $x\in X$. A subset $A\subseteq X$ is $G$-invariant or just invariant, if $gA=A$ for every $g\in G$, where $gA=\{gx:x\in A\}$. The topological dynamical system $(X,G)$ is minimal if for every $x\in X$ its orbit $O_{G}(x)=\{gx:g\in G\}$ is dense in $X$. A minimal component of $(X,G)$ is a closed non-empty $G$-invariant subset $Y$ of $X$ such that $(Y,G)$ is minimal. It is well known that every topological dynamical system has minimal components (see for example [2]). The topological dynamical system $(X,G)$ is equicontinuous if the homeomorphisms defined by the action of $G$, constitutes a uniformly equicontinuous family. When $X$ is a metric space, this is equivalent to say that for every $\varepsilon>0$ there exists $\delta>0$, such that if $d(x,y)<\delta$ then $d(gx,gy)<\varepsilon$, for every $x,y\in X$ and $g\in G$.
Let $(X,G)$ and $(Y,G)$ be two topological dynamical systems. We say that a map $\pi:X\to Y$ is equivariant if $\pi(gx)=g\pi(x)$ for every $g\in G$ and $x\in X$. The equivariant map $\pi$ is a factor map if it is also surjective. If $\pi:X\to Y$ is a factor map, we say that $(X,G)$ is a topological extension or an extension of $(Y,G)$, and $(Y,G)$ is a topological factor or a factor of $(X,G)$. If $\pi$ is bijective, we say that $(X,G)$ and $(Y,G)$ are conjugate. The factor map $\pi:X\to Y$ is almost one-to-one if the set $\{y\in Y:|\pi^{-1}\{\pi(y)\}|=1\}$ is dense in $Y$.
Every topological dynamical system $(X,G)$ has a unique (up to conjugacy) maximal equicontinuous factor, that is, a factor $(Y,G)$ which is equicontinuous, and such that every other factor from $(X,G)$ to an equicontinuous system, passes through the factor from $(X,G)$ to $(Y,G)$. A system $(X,G)$ is almost automorphic if its maximal equicontinuous factor is minimal, and if the associated factor map is almost one-to-one.
2.1. Invariant measures
Let $(X,G)$ be a topological dynamical system. We denote by ${\mathcal{M}}(X)$ the space of Borel probability measures of $X$. This space is compact, and metrizable if $X$ also is, with respect to the weak topology (see [9]). An invariant probability measure of $(X,G)$ is a measure $\mu\in{\mathcal{M}}(X)$ such that $\mu(gA)=\mu(A)$, for every $g\in G$ and every Borel set $A\subseteq X$. We denote ${\mathcal{M}}(X,G)$ the space of invariant probability measures of $(X,G)$. It is well known that the group $G$ is amenable if and only if for every topological dynamical system $(X,G)$ the space ${\mathcal{M}}(X,G)$ is non-empty (see, for example, [23]).
The next result is classical in Ergodic Theory (see [9, 26]).
Lemma 3.
Let $(X,G)$ and $(Y,G)$ be two topological dynamical systems. If $\pi:X\to Y$ is a factor map, then the function $\pi^{*}:{\mathcal{M}}(X)\to{\mathcal{M}}(Y)$ defined as
$\pi^{*}\mu(A)=\mu(\pi^{-1}(A)),\mbox{ for every }\mu\in{\mathcal{M}}(X)\mbox{ and every Borel set }A\subseteq Y,$
is affine, continuous and surjective.
In the context of Lemma 3, if the factor $(Y,G)$ has no invariant probability measures, then neither does the extension $(X,G)$. In other words, extensions of topological dynamical systems with no invariant measures, have no invariant measures.
2.2. Cantor systems and subshifts
We say that the topological dynamical system $(X,G)$ is a Cantor system if $X$ is a Cantor set, i.e., non-empty, compact, metric, totally disconnected and without isolated points. Examples of Cantor systems are given by the full $G$-shifts, which we recall in the next lines.
Let $\Sigma$ be a finite alphabet, i.e., a finite set with at least two elements. Consider the set
$\Sigma^{G}=\{x=(x(g))_{g\in G}:x(g)\in\Sigma,\mbox{ for every }g\in G\}.$
The space $\Sigma^{G}$ is a Cantor set when we endow $\Sigma$ with the discrete topology and $\Sigma^{G}$ with the product topology. The cylinder set given by the finite set $F\subseteq G$ and $\Lambda\in\Sigma^{F}$, is defined as
$[\Lambda,F]=\{x\in\Sigma^{G}:x(g)=\Lambda(g),\mbox{ for every }g\in F\}.$
The topology of $\Sigma^{G}$ is generated by the cylinder sets.
The group $G$ acts continuously on $\Sigma^{G}$ by means of the shift action, which is defined as follows: let $x=(x(g))_{g\in G}\in\Sigma^{G}$ and $h\in G$,
$hx(g)=x(h^{-1}g),\mbox{ for every }g\in G.$
The Cantor system $(\Sigma^{G},G)$ is called the $|\Sigma|$-full $G$-shift, where $|\Sigma|$ denotes the cardinality of $\Sigma$. A subshift of $\Sigma^{G}$ is a closed $G$-invariant subset $X\subseteq\Sigma^{G}$. We also call subshift the topological dynamical system $(X,G)$ given by the restriction of the shift-action of $G$ on $X$ (see [3] for more details on subshifts). The next lemma shows that minimal infinite subshifts are Cantor systems.
Lemma 4.
Let $(X,G)$ be a minimal topological dynamical system such that $X$ is totally disconnected. If $X$ is infinite, then $X$ has no isolated points.
Proof.
Suppose that $x\in X$ is an isolated point. Then $\{x\}$ is an open set of $X$. Since the system $(X,G)$ is minimal, the set of return times of $x$ to $\{x\}$ is syndetic (see [2]), which is equivalent to say that $\{g\in G:gx=x\}$ is a finite index subgroup of $G$. This implies that the orbit of $x$ is finite. Thus if $X$ is infinite, then $X$ has no isolated points. ∎
2.3. Residually finite groups and their Cantor systems.
The group $G$ is residually finite if for every $g\in G\setminus\{1_{G}\}$ there exist a finite group $F$ and a group homomorphism $\phi:G\to F$ such that $\phi(g)\neq 1_{F}$. This property is equivalent to the existence of a decreasing sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of finite index subgroups of $G$ such that $\bigcap_{n\in\mathbb{N}}\Gamma_{n}=\{1_{G}\}$. The residual subgroup $R(G)$ of $G$ is the intersection of all the finite index subgroups of $G$. The group $R(G)$ is normal in $G$ and coincides with the intersection of all the finite index normal subgroups of $G$. The group $G$ is residually finite if and only if $R(G)$ is trivial (see [3, Proposition 2.1.11]). Observe that $G/R(G)$ is a residually finite group, for every group $G$. Examples of residually finite groups are the free group $\mathbb{F}_{n}$ with $n$ generators and $\mathbb{Z}^{n}$, for every $n\in\mathbb{N}$.
2.4. Odometers
Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a strictly decreasing sequence of finite index subgroups of $G$. The $G$-odometer associated to $(\Gamma_{n})_{n\in\mathbb{N}}$ is defined as
[…]
where $\tau_{n}\colon G/\Gamma_{n+1}\to G/\Gamma_{n}$ is the canonical projection, for every $n\in\mathbb{N}$. The space $X$ is a Cantor set when we endow every $G/\Gamma_{n}$ with the discrete topology, $\prod_{n\in\mathbb{N}}G/\Gamma_{n}$ with the product topology and $X$ with the induced topology. Let $\tau:G\to X$ given by $\tau(g)=(g\Gamma_{n})_{n\in\mathbb{N}}$. The group $G$ acts on $X$ as $g(g_{n}\Gamma_{n})_{n\in\mathbb{N}}=(gg_{n}\Gamma_{n})_{n\in\mathbb{N}}$, for every $g\in G$ and $(g_{n}\Gamma_{n})_{n\in\mathbb{N}}\in X$. The topological dynamical system $(X,G)$ is known as the $G$-odometer or odometer associated to the sequence $(\Gamma_{n})_{n\in\mathbb{N}}$. This system is equicontinuous and minimal. The $G$-odometer is free if and only if $\bigcap_{n\in\mathbb{N}}g_{n}\Gamma_{n}g_{n}^{-1}=\{1_{G}\}$, for every $(g_{n}\Gamma_{n})_{n\in\mathbb{N}}\in\overleftarrow{G}$, which implies that there exists a free $G$-odometer if and only if $G$ is residually finite (see [6]).
If the groups $\Gamma_{n}$’s are normal, we call the $G$-odometer an exact $G$-odometer or exact-odometer, and we denote it as $\overleftarrow{G}=\overleftarrow{G}_{(\Gamma_{n})}$. In this case, $\overleftarrow{G}$ is a subgroup of $\prod_{n\in\mathbb{N}}G/\Gamma_{n}$, and the function $\tau:G\to\overleftarrow{G}$ is a group homomorphism. Furthermore, the $G$ action corresponds to the coordinatewise left multiplication by $\tau(g)$, for every $g\in G$. On the other hand, $(\overleftarrow{G},G)$ is free if and only if $\bigcap_{n\in\mathbb{N}}\Gamma_{n}=\{1_{G}\}$. In this case, $\tau(G)$ is a dense subgroup of $\overleftarrow{G}$ isomorphic to $G$, and $\overleftarrow{G}$ is a totally disconnected metric compactification of $G$. Indeed, the totally disconnected metric compactifications of $G$ are precisely the free exact $G$-odometers (see [4, Lemma 2.1]).
Proposition 2.
If $(X,G)$ is a minimal equicontinuous Cantor system, then $(X,G)$ is conjugate to a $G$-odometer.
Proof.
It is sufficient to follow the proof of [8, Theorem 2.7], and to note that if $(X,G)$ is not free, then the function $\varphi:G\to K$ is not necessarily injective. This implies that the action of $G$ on $K$ is not necessarily free. The intersection of the groups $G_{n}$’s mentioned there, is the stabilizer of the neutral element $e\in K$ (and maybe non-trivial). Then, the proof follows to conclude that $(X,G)$ is regularly recurrent, meaning that for every $x\in X$ and every clopen neighborhood $C$ of $x$, there exists a finite index subgroup $\Gamma$ of $G$ such that $\gamma x\in C$, for every $\gamma\in\Gamma$. Let $x\in X$ and $(U_{n})_{n\in\mathbb{N}}$ be a nested sequence of clopen sets such that $\bigcap_{n\in\mathbb{N}}U_{n}=\{x\}$. Then, by [6, Lemma 4], for every $n\in\mathbb{N}$ there exists a finite index subgroup $\Gamma_{n}$ such that $C_{n}=\overline{O_{\Gamma_{n}}(x)}\subseteq U_{n}$ is a clopen set verifying $C_{n}\cap\gamma C_{n}\neq\emptyset$ if and only if $\gamma C_{n}=C_{n}$, and if and only if $\gamma\in\Gamma_{n}$. The system $(X,G)$ is conjugate to the odometer given by the sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ (see the proof of [6, Theorem 2]). ∎
The following result is a consequence of Proposition 2
Proposition 3.
Let $G$ be a countable group. The following statements are equivalent:
  1. (1)
    There exists a minimal equicontinuous Cantor system $(X,G)$.
  2. (2)
    The group $G/R(G)$ is infinite, where $R(G)$ is the residual subgroup of $G$.
  3. (3)
    $G$ is an extension of an infinite residually finite group.
Proof.
Suppose that $(X,G)$ is a minimal equicontinuous system. By Proposition 2, the system $(X,G)$ is conjugate to a $G$-odometer. Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a decreasing sequence of finite index subgroups of $G$ such that $(X,G)$ is conjugate to the odometer associated to $(\Gamma_{n})_{n\in\mathbb{N}}$. Let $\Gamma=\bigcap_{n\in\mathbb{N}}\Gamma_{n}$. Since $X$ is a Cantor set (therefore, an infinite set), the groups $\Gamma_{n}$’s do not stabilize. This implies that $G/\Gamma$ is infinite. Since $R(G)\subseteq\Gamma$, we get that $G/R(G)$ is infinite. This shows (1) implies (2). Since $G/R(G)$ is residually finite, the implication (2) to (3) is direct from the exact sequence $1\to R(G)\to G\to G/R(G)\to 1$. Finally, if $G$ is an extension of an infinite residually finite group, then a normal subgroup $H$ of $G$ exists such that $G/H$ is infinite and residually finite. This implies there exists a strictly decreasing sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of finite index subgroup of $G/H$ such that $\bigcap\Gamma_{n}=\{1_{G/H}\}$. If $\pi:G\to G/H$ is the canonical projection, then $(\pi^{-1}(\Gamma_{n}))_{n\in\mathbb{N}}$ is a strictly decreasing sequence of finite index subgroups of $G$. The $G$-odometer associated to $(\pi^{-1}(\Gamma_{n}))_{n\in\mathbb{N}}$ is a minimal equicontinuous Cantor system given by the action of $G$. This shows (3) implies (1). ∎
2.5. Toeplitz subshifts
Let $\Sigma$ be a finite alphabet. An element $x\in\Sigma^{G}$ is Toeplitz if for every $g\in G$ there exists a finite index subgroup $\Gamma$ of $G$ such that $x(g)=x(\gamma^{-1}g)$, for every $\gamma\in\Gamma$. We say that $\Gamma$ is a period of $x$ or that the coordinate $g$ is periodic with respect to $\Gamma$. A subshift $X\subseteq\Sigma^{G}$ is a Toeplitz subshift if there exists a Toeplitz element $x\in\Sigma^{G}$, such that $X=\overline{O_{G}(x)}$. The Toeplitz subshifts are always minimal (see [6, 24, 25]). On the other hand, observe that if $x\in\Sigma^{G}$ is a non-periodic Toeplitz subshift (i.e., its stabilizer $\{g\in G:gx=x\}$ with respect to the shift action is trivial), then the intersection of its periods is trivial. This implies that $G$ is residually finite. In addition to the following proposition, this observation allows us to characterize residually finite groups as those that admit non-periodic Toeplitz elements.
Proposition 4.
([6, Theorem 7]) Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a strictly decreasing sequence of finite index subgroups (not necessarily normal or with trivial intersection) of $G$. Then there exists a Toeplitz subshift $X\subseteq\{0,1\}^{G}$ such that $(X,G)$ is an almost 1-1 extension of the $G$-odometer associated to $(\Gamma_{n})_{n\in\mathbb{N}}$.
2.5.1. Period structure of a Toeplitz element.
Let $x\in\Sigma^{G}$ be a Toeplitz element. Suppose that the finite index subgroup $\Gamma$ of $G$ is a period of $x$. We define
${\rm Per}(x,\Gamma,\alpha)$ =$=\{g\in G\colon x(\gamma g)=\alpha\mbox{ for every }\gamma\in\Gamma\},\mbox{ for every }\alpha\in\Sigma.$
${\rm Per}(x,\Gamma)$ $=\bigcup_{\alpha\in\Sigma}{\rm Per}(x,\Gamma,\alpha).$
The elements of ${\rm Per}(x,\Gamma)$ are those belonging to some coset $\Gamma g$ for which $x$ restricted to $\Gamma g$ is constant.
For every $g\in G$, for every $\alpha\in\Sigma$ and every period $\Gamma$ of $x$, we have $g{\rm Per}(x,g^{-1}\Gamma g,\alpha)={\rm Per}(gx,\Gamma,\alpha)$. The period $\Gamma$ of $x$ is an essential period of $x$ if ${\rm Per}(x,\Gamma,\alpha)\subseteq{\rm Per}(gx,\Gamma,\alpha)$ for every $\alpha\in\Sigma$, implies $g\in\Gamma$.
Lemma 5.
([4, Lemma 2.4], [6]) Let $x\in\Sigma^{G}$ be a Toeplitz element. For every period $\Gamma$ of $x$, there exists an essential period $K$ of $x$ such that ${\rm Per}(x,\Gamma,\alpha)\subseteq{\rm Per}(x,K,\alpha)$, for every $\alpha\in\Sigma$.
A period structure of $x$ is a nested sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of essential periods of $x$, such that $G=\bigcup_{n\in\mathbb{N}}\Gamma_{n}$. A period structure of $x$ always exists (see [6, Corollary 6]). Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a period structure of $x$. Let $X=\overline{O_{G}(x)}$ be the associated Toeplitz subshift. For each $n\in\mathbb{N}$, consider the set
$B_{n}=\{y\in X:{\rm Per}(y,\Gamma_{n},\alpha)={\rm Per}(x,\Gamma_{n},\alpha),\mbox{ for every }\alpha\in\Sigma\}.$
It is straightforward to verify that $gx\in B_{n}$ if and only if $g\in\Gamma_{n}$. This implies the following result.
Lemma 6.
([6, Lemma 8 and Proposition 6] or [4, Lemma 2.5]) For every $n\in\mathbb{N}$, the set $B_{n}$ is closed. In addition, for every $g,h\in G$ the following statements are equivalent:
  1. (1)
    $gB_{n}\cap hB_{n}\neq\emptyset.$
  2. (2)
    $gB_{n}=hB_{n}.$
  3. (3)
    $g\Gamma_{n}=h\Gamma_{n}$.
Then the collection $\{g^{-1}B_{n}:g\in D_{n}\}$ is a clopen partition of $X$, where $D_{n}$ is a subset of $G$ containing exactly one representative element of the classes in $G/\Gamma_{n}$, for every $n\in\mathbb{N}$.
Remark 7.
Observe that the stabilizer of $y\in X$ is the group $\bigcap_{n\in\mathbb{N}}v_{n}\Gamma_{n}v_{n}^{-1}$, where $v_{n}\in G$ is such that $y\in v_{n}B_{n}$. Thus, if the groups $\Gamma_{n}$ are normal subgroups of $G$, the Toeplitz subshift $(X,G)$ is free if and only if $\bigcap_{n\in\mathbb{N}}\Gamma_{n}=\{1_{G}\}$.
Proposition 5.
Let $x\in\Sigma^{G}$ be a Toeplitz element and let $X=\overline{O_{G}(x)}$. Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a period structure of $x$, and let $Y$ be the $G$-odometer associated to $(\Gamma_{n})_{n\in\mathbb{N}}$. The map $\pi:X\to Y$ given by
$\pi(x)=(g_{n}\Gamma_{n})_{n\in\mathbb{N}},\mbox{ where }x\in g_{n}B_{n},\mbox{for every }n\in\mathbb{N}$
is an almost 1-1 factor map. Since $(X,G)$ is minimal, this implies that $(Y,G)$ is the maximal equicontinuous factor of $X$.
Conversely, if $(Y,G)$ is a minimal equicontinuous Cantor system, then the symbolic almost 1-1 extensions of $(Y,G)$ are Toeplitz subshifts.
Proof.
The first part is direct from Lemma 6. The second part is a consequence of Proposition 2 and [25, Theorem 5.8]. ∎
3. Almost automorphic systems with no invariant measures.
Let $G$ be an infinite countable group, and let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a strictly decreasing sequence of finite index subgroups of $G$. Observe that $\bigcap_{n\in\mathbb{N}}\Gamma_{n}$ is not of finite index. The following result is well known; we include it for the reader’s convenience (see, for example, [4, 7]).
Lemma 8.
Let $G$ be an infinite countable group, and let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a strictly decreasing sequence of finite index subgroups of $G$. Let $H=\bigcap_{n\in\mathbb{N}}\Gamma_{n}$ and $D\subseteq G$ be a set that contains $1_{G}$ and exactly one element of each class $\{Hg:g\in G\}$. Then there exist an increasing sequence $(n_{i})_{i\in\mathbb{N}}\subseteq\mathbb{N}$ and a sequence $(D_{i})_{i\in\mathbb{N}}$ of finite subsets of $D$, such that for every $i\in\mathbb{N}$
  1. (1)
    $1_{G}\in D_{1}$ and $D_{i}\subseteq D_{i+1}$.
  2. (2)
    $D_{i}$ contains exactly one element of each class $\{\Gamma_{n_{i}}g:g\in G\}$.
  3. (3)
    $D=\bigcup_{i\in\mathbb{N}}D_{i}$.
Proof.
Observe that if $S\subseteq D$ is a finite set, then there exists $l\geq 1$ such that if $g_{1}$ and $g_{2}$ are two different elements in $S$, then $\Gamma_{l}g_{1}\neq\Gamma_{l}g_{2}$. Indeed, suppose there exists a finite subset $S$ of $D$ such that for every $l\geq 1$ there are two different elements $g_{l}$ and $h_{l}$ in $S$ such that $\Gamma_{l}g_{l}=\Gamma_{l}h_{l}$. Since $S$ is finite, there exist $g,h\in S$ and a subsequence $(\Gamma_{l_{i}})_{i\in\mathbb{N}}$ such that $g_{l_{i}}=g$ and $h_{l_{i}}=h$, for every $i\in\mathbb{N}$. This implies that $hg^{-1}\in\bigcap_{i\in\mathbb{N}}\Gamma_{l_{i}}=\bigcap_{i\in\mathbb{N}}\Gamma_{i}=H$, which is a contradiction with the fact that $g\neq h\in S\subseteq D$.
Let $D=\{g_{1},g_{2},\dots\}$, with $g_{1}=1_{G}$. We will construct a sequence $(F_{i})_{i\in\mathbb{N}}$ of finite subsets of $D$ verifying points (1), (2) and (3): choose $F_{1}$ as a fundamental domain of $G/\Gamma_{1}$ such that $g_{1}\in F_{1}$. We can choose $F_{1}$ contained in $D$ by replacing the elements $g$ of $F_{1}$ which are not in $D$, by elements in $D$ which are in $Hg$. We set ${n_{1}}=1$. Let $S_{2}=F_{1}\cup\{g_{1},g_{2}\}\subseteq D$. By the previous discussion, there exists $\Gamma_{n_{2}}\subseteq\Gamma_{n_{1}}$ satisfying $\Gamma_{n_{2}}g\neq\Gamma_{n_{2}}h$ for every pair of distinct elements $g,h\in S_{2}$. Thus, we can take $F_{2}\subseteq D$, fundamental domain of $G/\Gamma_{n_{2}}$, such that $S_{2}\subseteq F_{2}$. Suppose we have defined $\Gamma_{n_{k-1}}$ and $F_{k-1}\subseteq D$ such that $\Gamma_{n_{k-1}}\subseteq\Gamma_{n_{k-2}}$ and $F_{k-1}$ is a fundamental domain of $\Gamma_{n_{k-1}}$ containing the set $F_{k-2}\cup\{g_{1},\dots,g_{k-1}\}$. Now, consider $S_{k}=\{g_{1},g_{2},\dots,g_{k}\}\cup F_{k-1}\subseteq D$. By the previous discussion, we can choose $\Gamma_{n_{k}}\subseteq\Gamma_{n_{k-1}}$ such that the elements in $S_{k}$ are in different equivalence classes of $G/\Gamma_{n_{k}}$. Thus, we can get $F_{k}\subseteq D$ satisfying $S_{k}\subseteq F_{k}$.
Thus, we obtain (1) and (2). Additionally,
$D=\bigcup_{i=1}^{\infty}S_{i}\subseteq\bigcup_{i=1}^{\infty}F_{i}\subseteq D,$
which implies (3).
Let $(Z,G)$ be the $G$-odometer associated to the strictly decreasing sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of finite index subgroups of $G$ (see Section 2.4). Let $D$ be a fundamental domain of $\bigcap_{n\in\mathbb{N}}\Gamma_{n}$ containing $1_{G}$. Taking subsequences if necessary, Lemma 8 ensures the existence of a sequence $(D_{i})_{i\in\mathbb{N}}$ of finite subsets of $D$ such that for every $i\in\mathbb{N}$
  1. (1)
    $1_{G}\in D_{1}$ and $D_{i}\subseteq D_{i+1}$.
  2. (2)
    $D_{i}$ is a fundamental domain of $\Gamma_{i}$.
  3. (3)
    $D=\bigcup_{i\in\mathbb{N}}D_{i}$.
For every $n\in\mathbb{N}$, we define
$C_{n}=\{(g_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z:g_{n}\Gamma_{n}=1_{G}\Gamma_{n}\}.$
It is straightforward to check that the topology of $Z$ is generated by the collection of sets $\{gC_{n}:g\in D_{n},n\in\mathbb{N}\}$ (see for example, [6]). Let $\nu$ be the unique invariant probability measure of $(Z,G)$. Observe that
$\nu(gC_{n})=\frac{1}{|D_{n}|}=\frac{1}{[G:\Gamma_{n}]},\mbox{ for every }g\in G\mbox{ and }n\in\mathbb{N}.$
Let $S=(n_{i})_{i\in\mathbb{N}}$ be an increasing sequence of positive integers. We define
$X_{0}=X_{0}(S)=\{(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z:G=\bigcup_{i\in\mathbb{N}}t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}\}.$
This set is invariant with respect to the $G$ action on $Y$. Indeed, if $(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in X_{0}$ then given $h,g\in G$, there exists $i\in\mathbb{N}$ such that $g^{-1}h\in t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}$. This implies that $h\in(gt_{n_{i}})\Gamma_{n_{i}}D_{n_{i-1}}$, from which we get that $g(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in X_{0}$.
Lemma 9.
There exists a subsequence $(\Gamma_{n_{i}})_{i\in\mathbb{N}}$ such that $\nu(X_{0})=0$, where
$X_{0}=\{(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z:G=\bigcup_{i\in\mathbb{N}}t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}\}.$
Proof.
Take $n_{0}=1$. Suppose we have chosen $n_{0}<\cdots<n_{i-1}$. We take $n_{i}>n_{i-1}$ such that
$\frac{|D_{n_{i-1}}|}{|D_{n_{i}}|}<\frac{1}{2^{i+1}}.$
Let $(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z$ and $g\in G$ be such that there exists $j\in\mathbb{N}$ for which $g\in t_{n_{j}}\Gamma_{n_{j}}D_{n_{j-1}}$. This implies that $g^{-1}t_{n_{j}}\in D_{n_{j-1}}^{-1}\Gamma_{n_{j}}$. From this, we get
$g^{-1}(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in\bigcup_{j\in\mathbb{N}}\bigcup_{d\in D_{n_{j-1}}^{-1}}dC_{n_{j}}.$
This implies
$X_{0}\subseteq\bigcap_{g\in G}\bigcup_{j\in\mathbb{N}}\bigcup_{d\in D_{n_{j-1}}^{-1}}gdC_{n_{j}}.$
Let $A_{j}=\bigcup_{d\in D_{n_{j-1}}^{-1}}dC_{n_{j}}$. We have,
$\nu(gA_{j})=\frac{|D_{n_{j-1}}|}{|D_{n_{j}}|}<\frac{1}{2^{j+1}},\mbox{ for every }g\in G.$
Then
$\nu\left(\bigcup_{j\in\mathbb{N}}gA_{j}\right)\leq\frac{1}{2}\sum_{j\in\mathbb{N}}\frac{1}{2^{j}}=\frac{1}{2}.$
This implies
$\nu(X_{0})\leq\nu(\bigcap_{g\in G}\bigcup_{j\in\mathbb{N}}gA_{j})\leq\frac{1}{2},$
Since $X_{0}$ is invariant and $\nu$ is ergodic, we deduce that $\nu(X_{0})=0$. ∎
Proof of Theorem 1.
Let $(Z,G)$ be a minimal equicontinuous system, with $Z$ a Cantor set. By Proposition 2, $(Z,G)$ is conjugate to a $G$-odometer. Let $(\Gamma_{i})_{i\in\mathbb{N}}$ be the sequence of finite index subgroups that defines $(Z,G)$. By Lemma 8, we can assume that $(\Gamma_{i})_{i\in\mathbb{N}}$ is such that there exists a sequence $(D_{i})_{i\in\mathbb{N}}$ of finite subsets of $D$, fundamental domain of $\bigcap_{i\in\mathbb{N}}\Gamma_{i}=H$, such that for every $i\in\mathbb{N}$
  1. (1)
    $1_{G}\in D_{1}$ and $D_{i}\subseteq D_{i+1}$.
  2. (2)
    $D_{i}$ is a fundamental domain of $\Gamma_{i}$.
  3. (3)
    $D=\bigcup_{i\in\mathbb{N}}D_{i}$.
By Lemma 9, we can also assume that $\nu(X_{0})=0$, where
$X_{0}=\{(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z:G=\bigcup_{i\in\mathbb{N}}t_{i}\Gamma_{i}D_{i-1}\}.$
Let $(Y,G)$ be a minimal topological dynamical system, with $Y$ compact and metric, such that there exists a factor map $\pi:Y\to Z$.
For every $i\in\mathbb{N}$, let $C_{i}=\{(g_{k}\Gamma_{k})_{k\in\mathbb{N}}\in Z:g_{i}\Gamma_{i}=1_{G}\Gamma_{i}\}$. Observe that the collection $\{g\pi^{-1}C_{i}:g\in D_{i}\}$ is a clopen partition of $Y$. Furthermore, $g\pi^{-1}C_{i}\cap\pi^{-1}C_{i}\neq\emptyset$, if and only if $g\pi^{-1}C_{i}=\pi^{-1}C_{i}$, and if and only if $g\in\Gamma_{i}$.
For every $y\in Y$ and $i\in\mathbb{N}$, we define $t_{i}=t_{i,y}$ as the element in $D_{i}$ such that
$\pi(y)\in t_{i,y}C_{i}.$
Observe this is equivalent to say that $\pi(y)=(t_{i,y}\Gamma_{i})_{i\in\mathbb{N}}$. Thus, for every $g\in G$, we have
(3.1) $gt_{i,y}=t_{i,gy}\gamma,\mbox{ for some }\gamma\in\Gamma_{i}.$
We consider $Y^{G}$ equipped with the product topology and with the shift action, given by
$g(y(h))_{h\in G}=(y(g^{-1}h))_{h\in G},\mbox{ for every }g\in G\mbox{ and }(y(h))_{h\in G}\in Y^{G}.$
Fix $y_{0}\in\bigcap_{i\in\mathbb{N}}\pi^{-1}C_{i}$. We will define, inductively, a sequence $(\phi_{i}(y_{0}))_{i\in\mathbb{N}}$ of elements in $Y^{G}$ as follows:
$\phi_{1}(y_{0})(h)=\begin{cases}y_{0},&\mbox{ if }h\in\Gamma_{1},\\ h^{-1}y_{0},&\mbox{ otherwise.}\end{cases}$
For $i>1$,
$\phi_{i}(y_{0})(h)=\begin{cases}\phi_{i-1}(y_{0})(d),&\mbox{ if there exist }d\in D_{i-1}\mbox{ such that }h\in\Gamma_{i}d,\\ \phi_{i-1}(y_{0})(h),&\mbox{ otherwise.}\end{cases}$
Claim 1: For every $j\geq i$, we have $\phi_{j}(y_{0})|_{\Gamma_{i}D_{i-1}}=\phi_{i}(y_{0})|_{\Gamma_{i}D_{i-1}}$. More precisely,
$\phi_{j}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i}(y_{0})(d_{i-1}),\mbox{ for every }d_{i-1}\in D_{i-1}\mbox{ and }\gamma_{i}\in\Gamma_{i}.$
Proof of Claim 2: Let $d_{i-1}\in D_{i-1}$ and $\gamma_{i}\in\Gamma_{i}$. By definition of $\phi_{i}$, we have
(3.2) $\phi_{i}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i-1}(y_{0})(d_{i-1}).$
If $\gamma_{i}d_{i-1}\notin\Gamma_{i+1}D_{i}$, then
$\phi_{i+1}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i-1}(y_{0})(d_{i-1}).$
If $\gamma_{i}d_{i-1}\in\Gamma_{i+1}d_{i}$, for some $d_{i}\in D_{i}$, then
$\phi_{i+1}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i}(y_{0})(d_{i}).$
Since $d_{i}\in\Gamma_{i}d_{i-1}$, we have
$\phi_{i+1}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i}(y_{0})(d_{i})=\phi_{i-1}(y_{0})(d_{i-1})=\phi_{i}(y_{0})(\gamma_{i}d_{i-1}).$
This shows that $\phi_{i}(y_{0})|_{\Gamma_{i}D_{i-1}}=\phi_{i+1}(y_{0})|_{\Gamma_{i}D_{i-1}}$.
Let $j>i$ and suppose that $\phi_{k}(y_{0})|_{\Gamma_{i}D_{i-1}}=\phi_{i}(y_{0})|_{\Gamma_{i}D_{i-1}}$, for every $i\leq k\leq j$.
If $\gamma_{i}d_{i-1}\notin\Gamma_{j+1}D_{j}$, then
$\phi_{j+1}(y_{0})(\gamma_{i}d_{i-1})=\phi_{j}(y_{0})(\gamma_{i}d_{i-1}).$
By hypothesis, we get
$\phi_{j+1}(y_{0})(\gamma_{i}d_{i-1})=\phi_{j}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i}(y_{0})(\gamma_{i}d_{i-1}).$
If $\gamma_{i}d_{i-1}\in\Gamma_{j+1}d_{j}$, for some $d_{j}\in D_{j}$, then $d_{j}\in\Gamma_{i}d_{i-1}$, and
$\phi_{j+1}(y_{0})(\gamma_{i}d_{i-1})=\phi_{j}(y_{0})(d_{j})=\phi_{i}(y_{0})(d_{j}),$
by hypothesis. Since
$\phi_{i}(y_{0})(d_{j})=\phi_{i-1}(y_{0})(d_{i-1})=\phi_{i}(y_{0})(\gamma_{i}d_{i-1}),$
we get
$\phi_{j+1}(y_{0})(\gamma_{i}d_{i-1})=\phi_{i}(y_{0})(\gamma_{i}d_{i-1}).$
This proves Claim 1.
Claim 1 implies that $\lim_{j\to\infty}\phi_{j}(y_{0})(g)=\phi_{i}(y_{0})(g)$, for every $g\in\Gamma_{i}D_{i-1}$. Since $G=HD=\bigcup_{i\in\mathbb{N}}HD_{i-1}\subseteq\bigcup_{i\in\mathbb{N}}\Gamma_{i}D_{i-1}$, for every $g\in G$ there exists $i\in\mathbb{N}$ such that $g\in\Gamma_{i}D_{i-1}$. This implies that $\lim_{j\to\infty}\phi_{j}(y_{0})(g)$ exists, for every $g\in G$. We call $\phi(y_{0})$ the element in $Y^{G}$ such that $\phi(y_{0})(g)=\lim_{j\to\infty}\phi_{j}(y_{0})(g)$, for every $g\in G$.
Let define $\phi:Y\to Y^{G}$ as
$\phi(y)(g)=\left\{\begin{array}[]{ll}\phi(y_{0})(d_{i-1}),&\mbox{ if there exists }i\in\mathbb{N}\mbox{ and }d_{i-1}\in D_{i-1}\mbox{ such that }g\in t_{i,y}\Gamma_{i}d_{i-1},\\ g^{-1}y,&\mbox{ if }g\in G\setminus\bigcup_{i\in\mathbb{N}}t_{i,y}\Gamma_{i}D_{i-1}.\end{array}\right.,$
for every $y\in Y$. Observe that $\phi(y_{0})$ is the limit of the sequence $(\phi_{i}(y_{0}))_{i\in\mathbb{N}}$. On the other hand, $\phi(y)$ is well defined, for every $y\in Y$. Indeed, let $y\in Y$ and suppose that $g\in t_{i,y}\Gamma_{i}d_{i-1}\cap t_{j,y}\Gamma_{j}d_{j-1}$, for some $d_{i-1}\in D_{i-1}$ and $d_{j-1}\in D_{j-1}$, with $j>i$. This implies that $d_{j-1}\in\Gamma_{i}d_{i-1}$. Then $\phi(y_{0})(d_{j-1})=\phi_{i}(y_{0})(d_{j-1})=\phi_{i}(y_{0})(d_{i-1})=\phi(y_{0})(d_{i-1})$, which shows that $\phi(y)(g)$ is well defined.
Observe that $\phi(y)=\phi(y_{0})$, for every $y\in\bigcap_{i\in\mathbb{N}}\pi^{-1}C_{i}.$
Claim 2: $\phi$ is continuous and equivariant,
Proof of Claim 2: Let $y\in Y$ and $F\subseteq G$ be a finite set. Let $\varepsilon>0$. We can write $F=F_{1}\cup F_{2}$, where $F_{1}\subseteq\bigcup_{i\in\mathbb{N}}t_{i,y}\Gamma_{i}D_{i-1}$ and $F_{2}\subseteq G\setminus\bigcup_{i\in\mathbb{N}}t_{i,y}\Gamma_{i}D_{i-1}$. Since $F_{1}$ is finite, there exists $k\in\mathbb{N}$ such that $F_{1}\subseteq\bigcup_{i=1}^{k}t_{i,y}\Gamma_{i}D_{i-1}$. Consider the open set $U=\pi^{-1}(t_{k,y}C_{k})\cap B_{\delta}(y)$, where $\delta>0$ is such that if $y^{\prime}\in B_{\delta}(y)$, then $d(h^{-1}y,h^{-1}y^{\prime})<\varepsilon$, for every $h\in F$.
If $g\in F_{1}$, then there exists $1\leq l\leq k$ such that $g=t_{l,y}\gamma_{l}d_{l-1}$, for some $d_{l-1}\in D_{l-1}$ and $\gamma_{l}\in\Gamma_{l}$. Then, since $t_{l,y^{\prime}}=t_{l,y}$ for every $y^{\prime}\in U$, we have $\phi(y)(g)=\phi(y_{0})(d_{i-1})=\phi(y^{\prime})(g)$.
For $g\in F_{2}$, we have
$\phi(y)(g)=g^{-1}y\mbox{ and }\phi(y^{\prime})(g)=g^{-1}y^{\prime}.$
Thus, if $y^{\prime}\in U$ then $\phi(y^{\prime})\in\prod_{h\in F}B_{\varepsilon}(\phi(y)(h))\times\prod_{g\in G\setminus F}Y$. This shows that $\phi$ is continuous.
Let $g,h\in G$ and $i\in\mathbb{N}$ be such that $g^{-1}h\in\Gamma_{i}d_{i-1}$, for some $d_{i-1}\in D_{i-1}$. This implies that
$g\phi(y_{0})(h)=\phi(y_{0})(g^{-1}h)=\phi(y_{0})(d_{i-1}).$
On the other hand, $g^{-1}h\in\Gamma_{i}d_{i-1}$ implies that $h\in g\Gamma_{i}d_{i-1}=t_{i,gy_{0}}\Gamma_{i}d_{i-1}$, which implies that
$\phi(gy_{0})(h)=\phi(y_{0})(d_{i-1}).$
Combining the two previous equations, we get that
$\phi(gy_{0})=g\phi(y_{0}),\mbox{ for every }g\in G.$
The rest of the proof follows from minimality of $(Y,G)$ and the continuity of $\phi$. Indeed, if $y\in Y$ then there exists a sequence $(g_{i}y_{0})_{i\in\mathbb{N}}$ that converges to $y$. The continuity of $\phi$ implies that $(\phi(g_{i}y_{0}))_{i\in\mathbb{N}}$ converges to $\phi(y)$, and the continuity of the shift $G$-action on $Y^{G}$ implies that $(g\phi(g_{i}y_{0}))_{i\in\mathbb{N}}$ converges to $g\phi(y)$, for every $g\in G$. On the other hand, $g\phi(g_{i}y_{0})=\phi(gg_{i}y_{0})$, the continuity of $\phi$ and of the $G$-action on $Y$, imply that $(g\phi(g_{i}y_{0}))_{i\in\mathbb{N}}$ also converges to $\phi(gy)$. This shows that $\phi$ is equivariant.
Claims 2 implies that $\phi:Y\to\phi(Y)\subseteq Y^{G}$ is a topological factor map. Then, since $(Y,G)$ is minimal, we get the system system $(\phi(Y),G)$ is also minimal.
Let define $\psi:Y\to Z\times\phi(Y)$ by $\psi(y)=(\pi(y),\phi(y))$, for every $y\in Y$. Since $\psi$ is equivariant and continuous, the system $(\psi(Y),G)$ is minimal. The map $\tau:\psi(Y)\to Z$, given by the projection on the first coordinate, is almost one-to-one. Indeed, if $\tau(\pi(y),\phi(y))=\pi(y_{0})$, then $y\in\bigcap_{i\in\mathbb{N}}\pi^{-1}C_{i}.$ This implies that $\phi(y)=\phi(y_{0})$, and then $\tau^{-1}\{\pi(y_{0})\}=\{(\pi(y_{0}),\phi(y_{0}))\}$.
On the other hand, the map $\psi$ is one-to-one on $\pi^{-1}(Z\setminus X_{0})$. Indeed, suppose that $y\neq y^{\prime}\in\pi^{-1}(Z\setminus X_{0})$ are such that $\psi(y)=\psi(y^{\prime})$. Then $\pi(y)=\pi(y^{\prime})$. This implies that $g\notin\bigcup_{i\in\mathbb{N}}t_{i,y}\Gamma_{i}D_{i-1}$ if and only if $g\notin\bigcup_{i\in\mathbb{N}}t_{i,y^{\prime}}\Gamma_{i}D_{i-1}$. From this, we get $\phi(y)(g)=g^{-1}y$ and $\phi(y^{\prime})(g)=g^{-1}y^{\prime}$, which implies $\phi(y)\neq\phi(y^{\prime})$, leading to a contradiction. This shows that $\psi$ is one-to-one on $\pi^{-1}(Z\setminus X_{0})$.
Suppose that $\mu$ is an invariant probability measure of $(\psi(Y),G)$. By Lemma 9, this measure is supported on $\tau^{-1}(Z\setminus X_{0})$. On the other hand, if $(\pi(y),\phi(y))\in\tau^{-1}(Z\setminus X_{0})$ then $\pi(y)\in Z\setminus X_{0}$, which implies that $y\in\pi^{-1}(Z\setminus X_{0})$, and then $\tau^{-1}(Z\setminus X_{0})\subseteq\psi(\pi^{-1}(Z\setminus X_{0}))$. Thus, we can define
$\tilde{\mu}(A)=\mu(\psi(A\cap\pi^{-1}(Z\setminus X_{0})))\mbox{ for every Borel set }A\subseteq Y.$
By [20, Theorem 2.8], this is well defined. We have $\tilde{\mu}(Y)=1$. If $(A_{n})_{n\in\mathbb{N}}$ is a collection of disjoint Borel sets of $Y$, then the sets $\psi(A_{n}\cap\pi^{-1}(Z\setminus X_{0}))$ are disjoint. This implies that $\mu$ is a probability measure. The measure $\mu$ is $G$-invariant, because $\psi$ is equivariant and $\pi^{-1}(Z\setminus X_{0})$ is $G$-invariant. This shows that if $(Y,G)$ has no invariant probability measures, then $(\psi(Y),G)$ has no invariant probability measures either. Observe that the map $\mu\in{\mathcal{M}}(\psi(Y),G)\to\tilde{\mu}\in{\mathcal{M}}(Y,G)$ is the inverse of the affine continuous map $\psi^{*}:{\mathcal{M}}(Y,G)\to{\mathcal{M}}(\psi(Y),G)$, given by $\psi^{*}(\mu)=\mu\circ\psi^{-1}$, for every $\mu\in{\mathcal{M}}(Y,G)$. This shows that ${\mathcal{M}}(Y,G)$ and ${\mathcal{M}}(\psi(Y),G)$ are affine homeomorphic.
Finally, observe that if $Y$ is a Cantor set, then $\psi(Y)\subseteq Z\times Y^{G}$ is totally disconnected. Since $(\psi(Y),G)$ is minimal and $\psi(Y)$ is infinite (it is an extension of $Z$), then Lemma 4 implies that $\psi(Y)$ is a Cantor set. ∎
4. From general Cantor systems to subshifts.
The goal of this Section is to show Proposition 1, Theorem 2 and Corollary 1.
Lemma 10.
Let $(X,G)$ be a Cantor system with no invariant probability measures. Then there exists a subshift $(Z,G)$ with no invariant probability measures which is a topological factor of $(X,G)$.
Proof.
Let $({\mathcal{P}}_{n})_{n\in\mathbb{N}}$ be a nested sequence of finite clopen partitions of $X$ such that
$\{A\in{\mathcal{P}}_{n}:n\in\mathbb{N}\}$ spans the topology of $X$. For every clopen set $A\subseteq X$, let define
$Z_{A}=\left\{\mu\in{\mathcal{M}}(X):\mu(A)=\mu(gA)\mbox{ for every }g\in G\right\}.$
We claim there exists $n\in\mathbb{N}$ such that $\bigcap_{A\in{\mathcal{P}}_{n}}Z_{A}=\emptyset$. Suppose this is not true, then for every $n\in\mathbb{N}$ there exists $\mu_{n}\in\bigcap_{A\in{\mathcal{P}}_{n}}Z_{A}$. Let $(\mu_{n_{i}})_{i\in\mathbb{N}}$ be a subsequence that converges to $\mu\in{\mathcal{M}}(X)$. Given a clopen set $A\subseteq X$ there exists $i_{0}\in\mathbb{N}$ such that for every $i\geq i_{0}$, there exist $A_{i,1},\cdots,A_{i,k_{i}}$ in ${\mathcal{P}}_{n_{i}}$ such that $A=\bigcup_{j=1}^{k_{i}}A_{i,j}$. This implies that for every $i\geq i_{0}$, and every $g\in G$,
$\mu_{n_{i}}(A)=\sum_{j=1}^{k_{i}}\mu_{n_{i}}(A_{i,j})=\sum_{j=1}^{k_{i}}\mu_{n_{i}}(gA_{i,j})=\mu_{n_{i}}(gA).$
Since $A$ and $gA$ are clopen, we have $\mu(A)=\lim_{i\to\infty}\mu_{n_{i}}(A)$ and $\mu(gA)=\lim_{i\to\infty}\mu_{n_{i}}(gA)$, which implies that $\mu$ is invariant, contradicting that ${\mathcal{M}}(X,G)=\emptyset$. This shows the claim.
Let $n\in\mathbb{N}$ be such that $\bigcap_{A\in{\mathcal{P}}_{n}}Z_{A}=\emptyset$. Let ${\mathcal{P}}_{n}=\{A_{1},\cdots,A_{k}\}$ and $\Sigma=\{1,\cdots,k\}$. Let $\pi:X\to\Sigma^{G}$ be the function defined as
$\pi(x)=(x(g))_{g\in G},\mbox{ where }x(g)=i\mbox{ if and only if }g^{-1}x\in A_{i}.$
The map $\pi$ is continuous and equivariant with respect to the shift action on $\Sigma^{G}$. This implies that $\pi:X\to\pi(X)$ is a factor map. The set $Z=\pi(X)$ is a subshift of $\Sigma^{G}$. The factor map $\pi$ induces an affine surjective and continuous transformation $\pi^{*}:{\mathcal{M}}(X)\to{\mathcal{M}}(Z)$ given by $\pi^{*}\mu(A)=\mu(\pi^{-1}(A))$, for every Borel set $A\subseteq Z$ (see Lemma 3). Suppose there exists $\mu_{0}\in{\mathcal{M}}(Z,G)$ and let $\mu\in(\pi^{*})^{-1}\{\mu_{0}\}$. For every $1\leq i\leq k$ let $C_{i}=\{x\in Z:x(1_{G})=i\}$. For every $g\in G$ we have $gA_{i}=g\pi^{-1}C_{i}=\pi^{-1}gC_{i}$, which implies
$\mu(gA_{i})=\mu(\pi^{-1}gC_{i})=\mu_{0}(gC_{i})=\mu_{0}(C_{i})=\mu(\pi^{-1}C_{i})=\mu(A_{i}).$
From this, we get that $\mu\in\bigcap_{A\in{\mathcal{P}}_{n}}Z_{A}$, which is a contradiction. This shows that ${\mathcal{M}}(Z,G)$ has no invariant probability measures. ∎
The proof of Theorem 1, will be the consequence of a series of results that we will show below.
Lemma 11.
Let $G$ be a non-amenable countable group. Let $(Y,G)$ be a minimal topological system verifying the following properties:
  1. (1)
    There exist a minimal Cantor system $(X,G)$, with no invariant probability measures, and an almost one-to-one factor map $\pi_{X}:X\to Y$.
  2. (2)
    There exist a minimal subshift $(Z,G)$, and factor maps $\pi_{Z}:Z\to Y$ and $\pi:X\to Z$.
  3. (3)
    $\pi_{X}=\pi_{Z}\circ\pi$.
Then there exists a minimal subshift $(\tilde{Z},G)$, with no invariant probability measures, that is an almost 1-1 extension of $(Y,G)$.
Proof.
From Lemma 10, there exist a minimal subshift $(W,G)$, with no invariant probability measures, and a topological factor map $\phi:X\to W$. Let $\tilde{\pi}:X\to W\times Z$ be the map given by $\tilde{\pi}(x)=(\phi(x),\pi(x))$, for every $x\in X$. The map $\phi$ is continuous and equivariant, which implies that $\tilde{\pi}:X\to\tilde{Z}=\tilde{\pi}(X)\subseteq W\times Z$ is a factor map. Observe that $(\tilde{Z},G)$ is a minimal subshift of $(\Sigma_{W}\times\Sigma_{Z})^{G}$, where $\Sigma_{W}$ and $\Sigma_{Z}$ are the finite alphabets on which $W$ and $Z$ are defined, respectively. Let $\tau_{W}:\tilde{Z}\to W$ and $\tau_{Z}:\tilde{Z}\to Z$ the respective projection maps. These maps are continuous, equivariant, and by the minimality of $W$ and $Z$, they satisfy $\tau_{W}(\tilde{Z})=W$ and $\tau_{Z}(\tilde{Z})=Z$. Since $(\tilde{Z},G)$ is an extension of $(W,G)$, the space of invariant probability measures of $(\tilde{Z},G)$ is empty. On the other hand, $\tau:\tilde{Z}\to Y$ defined by $\tau=\pi_{Z}\circ\tau_{Z}$, is a factor map. By hypothesis, there exists $y_{0}\in Y$ such that $|\pi_{X}^{-1}\{y_{0}\}|=1$. Let $z_{1}$ and $z_{2}$ in $\tilde{Z}$ such that $\tau(z_{1})=\tau(z_{2})=y_{0}$. Let $x_{1},x_{2}\in X$ be such that $z_{1}=(\phi(x_{1}),\pi(x_{1}))$ and $z_{2}=(\phi(x_{2}),\pi(x_{2}))$. We have $\tau(z_{1})=\pi_{Z}(\pi(x_{1}))=\pi_{X}(x_{1})=y_{0}$, which implies that $x_{1}\in\pi_{X}^{-1}\{y_{0}\}$. Similarly, we get $x_{2}\in\pi_{X}^{-1}\{y_{0}\}$, which implies $x_{1}=x_{2}$ and then $z_{1}=z_{2}$. This shows that $(\tilde{Z},G)$ is an almost 1-1 extension of $(Y,G)$.
Figure 1. Diagram corresponding to the different maps of Lemma 11.
The following lemma is a well known result. Nevertheless, we add a proof for completeness.
Lemma 12.
Let $(X,G),(Y,G)$ and $(Z,G)$ be topological dynamical systems. Suppose that $(Y,G)$ minimal. If $\pi:X\to Y$ and $\phi:Y\to Z$ are two almost $1$-$1$ factor maps, then $\phi\circ\pi:X\to Z$ is an almost $1$-$1$ factor map.
Proof.
Let $S=\{y\in Y:|\pi^{-1}(y)|=1\}$ and $T=\{z\in Z:|\phi^{-1}(z)|=1\}$ be the residual sets given by the almost $1$-$1$ factor maps $\pi$ and $\phi$, respectively. Consider $y\in\phi^{-1}(T)$. Since $T$ is $G$-invariant and $\phi$ is a factor map, the orbit of $y$ is contained in $\phi^{-1}(T)$. Thus the minimality of $Y$ implies that $\phi^{-1}(T)$ is a $G_{\delta}$ dense set in $Y$. By Baire’s category theorem, we have $\phi^{-1}(T)\cap S\neq\emptyset$. Take $y\in\phi^{-1}(T)\cap S$. Let $x$ and $x^{\prime}$ be two elements of $X$ such that $\phi(\pi(x))=\phi(\pi(x^{\prime}))=\phi(y)$. Since $\phi(y)\in T$, we have $\pi(x)=\pi(x^{\prime})=y\in S$. Consequently, $x=x^{\prime}$, and since $Z$ is minimal, we conclude that $\phi\circ\pi$ is an almost $1$-$1$ factor map. ∎
Lemma 13.
Let $(X,G)$ and $(Y,G)$ be two topological dynamical systems. Suppose that $\pi:X\to Y$ is an almost 1-1 factor map. Let $T\subseteq\{y\in Y:|\pi^{-1}(y)|=1\}$. Then the map $\hat{\pi}:T\to X$, given by $\hat{\pi}(y)=z$, where $\{z\}=\pi^{-1}\{y\}$, is continuous, injective and $\hat{\pi}^{-1}|_{\hat{\pi}(T)}$ is continuous.
Proof.
Let $C\subseteq X$ be a closed set. Since $\pi$ is a continuous map between compact Hausdorff spaces, the set $\pi(C)$ is closed. Let us show that $\pi(C)\cap T=\hat{\pi}^{-1}(C)$. If $y\in\pi(C)\cap T$, then there exists a unique $z\in\pi^{-1}\{y\}$ with $z\in C$. Since $\hat{\pi}(y)=z\in C$, we get $\pi(C)\cap T\subseteq\hat{\pi}^{-1}(C)$. On the other hand, if $y\in\hat{\pi}^{-1}(C)$, then $y\in T$ and $\hat{\pi}(y)=\pi^{-1}(y)=z\in C$, which implies that $y\in\pi(C)\cap T\subseteq\pi(C)$. Thus we conclude that $\pi(C)\cap T=\hat{\pi}^{-1}(C)$. This shows that the pre-image by $\hat{\pi}$ of every closed set is closed in $T$, which implies that $\hat{\pi}$ is continuous. The injectivity follows from the fact that $T\subseteq\{y\in Y:|\pi^{-1}(y)|=1\}$. Since $\hat{\pi}^{-1}|_{\hat{\pi}(T)}=\pi|_{\hat{\pi}(T)}$, the map $\hat{\pi}^{-1}|_{\hat{\pi}(T)}$ is continuous. ∎
Lemma 14.
Let $(Y,G)$ be a minimal topological dynamical system verifying the following properties:
  1. (1)
    There exist a minimal Cantor system $(X,G)$ and an almost one-to-one factor map $\phi:X\to Y$.
  2. (2)
    There exist a subshift $(Z,G)$ and an almost one-to-one factor map $\pi_{Z}:Z\to Y$.
Then there exist
  • A minimal Cantor system $(\tilde{X},G)$ which is a topological extension of $(X,G)$.
  • A factor map $\pi:\tilde{X}\to Z$ and an almost one-to-one factor map $\pi_{\tilde{X}}:\tilde{X}\to Y$
such that
$\pi_{\tilde{X}}=\pi_{Z}\circ\pi.$
Proof.
Denote $\mathcal{T}=\{y\in Y:|\pi_{Z}^{-1}\{y\}|=1\}$, and $H=\{y\in Y:|\phi^{-1}\{y\}|=1\}$ the $G$-invariant residual sets given by the almost $1$-$1$ factor maps $\pi_{Z}$ and $\phi$, respectively. Note that $T=\mathcal{T}\cap H$ is still a $G$-invariant residual set, thanks to the Baire’s category theorem, and consequently, a non-empty set. Consider the maps $\hat{\pi_{Z}}:T\to Z$ and $\hat{\phi}:T\to X$, such that $\hat{\pi_{Z}}(y)$ and $\hat{\phi}(y)$ are the unique elements in $\pi_{Z}^{-1}\{y\}$ and $\phi^{-1}\{y\}$, respectively, for every $y\in T$. Lemma 13 implies that these maps are continuous, injective with continuous inverse maps. Denote $X^{\prime}=\phi^{-1}(T)$, and let define $\tilde{X}=\overline{\{(x,\hat{\pi_{Z}}(\phi(x))):x\in X^{\prime}\}}\subset X\times Z$. The group $G$ acts on $X\times Z$ coordinatewise. The set $\tilde{X}$ is invariant with respect to this action, because is the closure of an invariant set. Let $\tau_{X}:\tilde{X}\to X$ and $\pi:\tilde{X}\to Z$ the respective projection maps. These maps are equivariant, and the minimality of $X$ and $Z$ imply that $\tau_{X}$ and $\pi$ are surjective, then $\tau_{X}$ and $\pi$ are factor maps. Moreover, $\pi$ is almost $1$-$1$. Indeed, let $x\in X^{\prime}$ and suppose that $(x^{\prime},y^{\prime})\in\tilde{X}$ is such that $\pi(x^{\prime},y^{\prime})=\pi(x,\hat{\pi_{Z}}(\phi(x)))$. This implies that $y^{\prime}=\hat{\pi_{Z}}(\phi(x))$. Suppose that $x^{\prime}\neq x$. Since $X$ is Hausdorff, there exist disjoint open neighborhoods $U\subseteq X$ and $V\subseteq X$ of $x$ and $x^{\prime}$, respectively. The continuity of $\hat{\phi}$ and $\pi_{Z}$ imply that $\pi_{Z}^{-1}(\hat{\phi}^{-1}(U))$ is an open set of $Z$. Moreover, since $x\in U\cap X^{\prime}$, we conclude that $\phi(x)\in\hat{\phi}^{-1}(U)\cap T$. Thus, we have that $\hat{\pi_{Z}}(\phi(x))\in\pi_{Z}^{-1}(\phi^{-1}(U))$. Observe that $V\times\pi_{Z}^{-1}(\hat{\phi}^{-1}(U))$ is an open neighborhood of $(x^{\prime},\hat{\pi_{Z}}(\phi(x)))$. Therefore, there exists $w\in X^{\prime}$ such that $(w,\hat{\pi_{Z}}(\phi(w)))\in V\times\pi_{Z}^{-1}(\hat{\phi}^{-1}(U))$. Using that $\hat{\pi_{Z}}(\phi(w))\in\pi_{Z}^{-1}(\hat{\phi}^{-1}(U))$, we have $\phi(w)\in\hat{\phi}^{-1}(U)\cap T$ and consequently, $w\in U$. Hence, we obtain a contradiction with the fact that $U\cap V=\emptyset$. Thus, $x=x^{\prime}$ and $\pi$ is an almost 1-1 factor.
Since $\pi_{Z}$ and $\pi$ are almost 1-1 factor maps, Lemma 12 implies that $\pi_{\tilde{X}}=\pi_{Z}\circ\pi$ is an almost 1-1 factor map. Finally, we will show that $\tilde{X}$ is minimal. Since the image by $\pi$ of every closed $G$-invariant set in $\tilde{X}$ is a closed $G$-invariant set in $Z$, the minimality of $Z$ implies that for every $\tilde{x}\in\tilde{X}$, we have $\pi(\overline{\{g\tilde{x}:g\in G\}})=Z$. This implies that $\{\hat{\pi_{Z}}(\phi(x)):x\in X^{\prime}\}\subseteq\pi(\overline{\{g\tilde{x}:g\in G\}})$. On the other hand, previously we have shown that every element in $\{\hat{\pi_{Z}}(\phi(x)):x\in X^{\prime}\}$ has a unique pre-image by $\pi$, which implies that
$\{(x,\hat{\pi_{Z}}(\phi(x))):x\in X^{\prime}\}\subseteq\overline{\{g\tilde{x}:g\in G\}}$. Since $\{(x,\hat{\pi_{Z}}(\phi(x))):x\in X^{\prime}\}$ is dense in $\tilde{X}$, we get that $\overline{\{g\tilde{x}:g\in G\}}=\tilde{X}$. Since $\tilde{X}$ is totally disconnected (it is a subspace of $X\times Z$) and infinite (it is an extension of $X$), by Lemma 4 we get that $(\tilde{X},G)$ is a minimal Cantor system.
Figure 2. Diagram corresponding to the different maps of Lemma 14.
Proof of Proposition 1.
Let $(Y,G)$, $(X,G)$ and $(Z,G)$ be the topological dynamical systems of Proposition 1. Applying Lemma 14 to these three systems, we get a minimal Cantor system $(\tilde{X},G)$ that is an extension of $(X,G)$. Since $(X,G)$ has no invariant probability measures, neither does $(\tilde{X},G)$. In addition, we get a factor map $\pi:\tilde{X}\to Z$, and an almost one-to-one factor map $\pi_{\tilde{X}}:\tilde{X}\to Y$, such that $\pi_{\tilde{X}}=\pi_{Z}\circ\pi.$
Now, applying Lemma 11 to $(Y,G)$, $(\tilde{X},G)$, and $(Z,G)$, we get there exists a minimal subshift $(\tilde{Z},G)$, with no invariant measures, which is an almost 1-1 factor of $(Y,G)$. ∎
Proof of Theorem 2.
Let $G$ be a non-amenable group. Let $(Z,G)$ be a minimal equicontinuous Cantor system. Lemma 2 implies that $(Z,G)$ is conjugate to the odometer associated to a strictly decreasing sequence of finite index subgroups $(\Gamma_{n})_{n\in\mathbb{N}}$. Since $G$ is non-amenable, there exists a minimal Cantor system $(\tilde{Y},G)$ with no invariant probability measures (see [18]). Consider the system $(Y,G)$, where $Y=\tilde{Y}\times Z$ equipped with the product action. Let $W\subseteq Y$ be a minimal component of $(Y,G)$. Since $(W,G)$ is an extension of $(\tilde{Y},G)$, we have that $(W,G)$ is a minimal Cantor system with no invariant probability measures. Since $(W,G)$ is also an extension of $(Z,G)$, applying Theorem 1, we get there exists a minimal Cantor system $(X,G)$ with no invariant probability measures, which is an almost one-to-one extension of $(Z,G)$. By 4, $(Z,G)$ admits symbolic almost one-to-one extension. Thus, applying Proposition 1, we get there exists a minimal subshift $(\tilde{Z},G)$, with no invariant probability measures, which is an almost one-to-one extension of $(Z,G)$. Finally, from Proposition 5 we deduce that $(\tilde{Z},G)$ is a Toeplitz $G$-subshift. ∎
Proof of Corollary 1.
The implications (1) to (2), (2) to (3), (3) to (4), and (4) to (5) are direct. The implication (4) to (1) is consequence of Theorem 2 and Proposition 3. ∎